# Scalable LLM Math Reasoning Acceleration with Low-rank Distillation

### v2

@@ Abstract
Due to long generations, large language model (LLM) math reasoning demands significant computational resources and time. While many existing efficient inference methods have been developed with excellent performance preservation on language tasks, they often severely degrade math performance. In this paper, we propose Caprese, a resource-efficient distillation method to recover lost capabilities from deploying efficient inference methods, focused primarily in feedforward blocks. With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the math capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs. Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>16% time-to-next-token reduction) while encouraging response brevity (up to 8.5% fewer tokens).
@@ 1 Introduction
With the increasing capabilities of large language models (LLMs) [VSP+17], evaluations are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving. These tasks tend to demand long generation which drives up latency, making efficiency a dire issue. Fortunately, many sparsity-based efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization. However, for math reasoning tasks, many of these algorithms begin to break down, decimating performance, despite their robustness in language settings. Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and math capabilities.
@@
One of the main differences between math reasoning and many language tasks is the generation length. Reasoning typically involves long generations from chain-of-thoughts (CoTs) [WWS+22], which often far surpass the length of the input query. However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time. For instance, deploying CATS [LLZ+24], a sparse thresholding method, on Gemma 2 2B [TRP+24] has little impact on language generation performance, but knocks GSM8K [CKB+21] accuracy from 51.02% to 34.42% (Table 1). Similarly, for thinking models: applying GRIFFIN [DCC24], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [GYZ+25] drops MATH-500 [LKB+23] accuracy from 79.40% to 42.00% (Table 2). Moreover, since generation is much more computationally demanding than prefill, there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling. Thus, this poses two key inference challenges with math reasoning. First, even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24]. Second, the already computationally expensive LLM autoregressive decoding process is accentuated with generating CoTs. This problem is further exaggerated as CoTs scale in length in thinking models [GYZ+25] and as the number of repeated generations scale to elicit higher quality answers [BJE+24, WSL+24, SLXK24]. An ideal method should be efficient, performance preserving, and easy to integrate.
@@
Figure 1: A full FF block maximizes accuracy without any benefit to efficiency. Sparse FF algorithms can be very efficient by using subsets of the FF block but harm math performance. Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for performative inference in language and math settings while being efficient. Layers are drawn as trapezoids to highlight the relative size of inputs and outputs.
@@
Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible. (We choose to focus on FF blocks since they contribute around 2/3 of an LLM’s parameters and about 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods. Using an oracle top-$k$ filter on FF nonlinearity output magnitudes, we observe huge reductions in error with a low-rank approximation to the FF output residuals (Figure 2). Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively tiny. This observation motivates us to estimate the residual from sparse FF methods with low-rank layers.
@@
We introduce Caprese (CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1). While Caprese works for any task, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities. Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome the aforementioned challenges of math reasoning and makes significant progress towards an ideal efficient method:
@@
    Performance Enhancement: Demonstrated across a variety of instruct and thinking models, Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks. Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling.
@@
    Low Budget: The distillation process of Caprese is cheap since we are able to see significant gains in math performance by training on only 20K synthetic math samples. Moreover, with a low-rank layer size of only 256, this equates to adding roughly $0.8\%$ additional parameters into Llama 3.1 8B and Gemma 2 9B, dwarfed by savings in active parameters ($\sim$2B cut for both models with Caprese (CATS)).
@@
Section 2 covers background on related works and describes GRIFFIN and CATS, two efficient sparse FF algorithms that serve as baselines. Section 3 details Caprese’s architecture and distillation procedure. Then, we showcase the strong performance of Caprese in Section 4: instruct LLM inference (Section 4.1), scaling generation outputs (Section 4.2), and scaling generation length with thinking LLMs (Section 4.3). Finally, we quantify efficiency improvements in Section 4.4.
@@ 2 Background
In this section, we provide an overview of the efficient LLM inference literature that Caprese builds upon with in-depth descriptions of GRIFFIN and CATS which we use as baselines and some recent developments for test-time scaling.
@@ 2.1 Related Works
To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23]. Pruning sets a parameter subset to zero to enforce static sparsity [LDS89, SLBK23, FA23, MFW23]. Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [JJNH91, ZLL+21, DDZ+24, LWD+23]. Similar to MoEs, another line of work aims at exploiting contextual sparsity to dynamically select FF neurons for each input without training. GRIFFIN [DCC24] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation. CATS [LLZ+24] uses hard thresholding to skip computation in parts of the FF block with a custom kernel.
@@
Several works aim to improve reasoning efficiency by directly reducing the number of generated tokens [RG24, NRS+24, AZ25, ABH25, XXZH25, QYS+25]. Since we aim to reduce the per-token latency, this line of research to encourage brevity is orthogonal to ours and could be used alongside our method.
@@
To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [KMH+20, HBM+22]. More recently, there is an increasing effort towards test-time scaling to boost performance on difficult tasks like math. Two ways of test-time scaling that have seen great success. First, for a single prompt, multiple responses can be sampled from the LLM [BJE+24, SLXK24, WSL+24, MSE24, SHZ+24]. This increases the probability that a desired response lies in the pool of responses. While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one. Second, CoTs can be lengthened for each response [GYZ+25, MYS+25, YLY+25]. By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
2.2 Training-free & Adaptively Sparse FF Methods
@@
where $\bm{W}_{g},\bm{W}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$, $\bm{W}_{2}\in\mathbb{R}^{D_{\text{FF}}\times D}$, $\sigma$ is a nonlinear function, $\odot$ is an element-wise multiplication operator, and $D_{\text{FF}}$ is the FF intermediate feature size. Typically, $D_{\text{FF}}\gg D$. Although recent LLMs use this architecture, for LLMs without gated functions (e.g., OPT [ZRG+22]), $\bm{X}\bm{W}_{1}$ can be removed. Bias terms are omitted for brevity.
@@
GRIFFIN [DCC24] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns. GRIFFIN calculates $\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$. The result $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ gives a metric to perform top-$k$ selection across corresponding columns of $\bm{W}_{g}$ and $\bm{W}_{1}$, and rows of $\bm{W}_{2}$ to produce $\widehat{\bm{W}}_{g},\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{k\times D}$. Then, for the generation phase, the following FF block is used for input $\bm{x}\in\mathbb{R}^{D}$:
@@
CATS [LLZ+24] uses hard thresholding to skip computation in part of the FF block. Letting $T_{\tau}$ be the hard thresholding function with threshold $\tau$, CATS computes
@@ 3.1 Layer-wise Distillation
Inference efficiency algorithms often introduce feature approximation errors in favor of faster generation, which we mitigate with distillation. Let the efficient and approximate FF block be $\widehat{\text{FF}}(\bm{x})=\widehat{\text{FF}}_{2}(\widehat{\text{FF}}_{1}(\bm{x}))$. In our design of Caprese, we do not constrain $\widehat{\text{FF}}$ to be any specific method or architecture. For instance, $\widehat{\text{FF}}$ could be an FF block with GRIFFIN [DCC24] or CATS [LLZ+24]. Then, the error is
@@
for input set $\mathcal{X}$, $\bm{L}\in\mathbb{R}^{D\times r}$, and $\bm{R}\in\mathbb{R}^{r\times D}$ where $r\ll D_{\text{FF}}$. The optimal solution can be computed analytically since this is a reduced rank regression problem, but the size of $|\mathcal{X}|$ and $D$ may make it prohibitively expensive. Therefore, we opt to learn $\bm{L}$ and $\bm{R}$ independently for every FF block (i.e., previous FF blocks are assumed to be from the original model), allowing for parallel layer-wise training. This takes inspiration from LESS [DYZ+24] which uses layer-wise training on attention residuals for key-value cache compression. Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen. We also distill end-to-end (E2E) to further improve the performance.
@@ 3.3 Parallel Inference Computation
The computation of $\bm{x}\bm{L}\bm{R}$ can be done in parallel with the original FF operations. In fact, $\bm{L}$ and $\bm{R}$ can be concatenated with the up and down projection matrices, respectively. In other words, the Caprese FF block is $\widehat{\text{FF}}^{+}(\bm{x})=\widehat{\text{FF}}^{+}_{2}(\widehat{\text{FF}}^{+}_{1}(\bm{x}))$, such that
@@ 3.4 Training Details
We set the inner dimension of our low-rank layer to $r=256$ (to see the effect of different $r$, see Appendix A). In comparison to the large inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B). We use a 20K subset of a common synthetic math training set for training. For a fair comparison, the same subset is used for both layer-wise and E2E distillation for every model. Each training sample is prepended with a CoT instruction: “Please reason step by step.” At test time, the actual instructions may be vastly different. Layer-wise and E2E training consists of 20 epochs and 3 epochs, respectively. Training and inference are done in BF16. Hyperparameters are listed in Appendix E.
@@ 3.5 Comparison with LoRA
Our method derives inspiration from and shares a slight connection with low-rank adaptation (LoRA), a widely used method for efficient fine tuning with the addition of low-rank parameters [HSW+22], but remains distinct since they target different inefficiencies. To illustrate, recall a sparse algorithm constructs $\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$ by selecting columns from $\bm{W}_{1}$. Alternatively, construct $\widetilde{\bm{W}}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$ by setting unwanted columns to zero, so nonzero columns in $\widetilde{\bm{W}}_{1}$ match columns in $\widehat{\bm{W}}_{1}$ and vice versa. Doing the same to columns of $\bm{W}_{g}$ and rows of $\bm{W}_{2}$, LoRA learns $\bm{A}_{1},\bm{A}_{g},\bm{B}_{2}^{\top}\in\mathbb{R}^{D\times r}$ and $\bm{B}_{1},\bm{B}_{g},\bm{A}_{2}^{\top}\in\mathbb{R}^{r\times D_{\text{FF}}}$ to form a new FF block with LoRA weights:
@@
Model GSM8K MATH CoQA QASPER
Llama 3.1 8B Instruct 27.90 13.94 63.88 15.16
GRIFFIN 17.44 6.16 63.37 12.63
Layer-wise Caprese 27.14 9.72 65.05 14.21
E2E Caprese 40.49 12.64 65.50 15.35
CATS 40.56 11.80 58.85 12.26
Layer-wise Caprese 37.68 13.66 63.92 14.34
E2E Caprese 51.86 13.88 64.27 14.42
Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
GRIFFIN 7.13 5.42 56.05 14.11
Layer-wise Caprese 13.72 6.62 56.07 13.40
E2E Caprese 21.00 8.44 56.55 13.88
CATS 19.18 7.54 54.40 13.88
Layer-wise Caprese 18.65 8.28 55.58 14.35
E2E Caprese 20.39 9.04 56.12 13.75
Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
GRIFFIN 28.96 10.98 64.52 12.52
Layer-wise Caprese 40.18 13.70 64.33 11.60
E2E Caprese 44.66 16.96 64.83 12.35
CATS 41.24 12.04 58.87 11.36
Layer-wise Caprese 45.49 13.62 60.53 12.26
E2E Caprese 46.85 14.40 61.72 12.36
Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
GRIFFIN 33.74 11.32 63.28 11.07
Layer-wise Caprese 42.53 12.32 63.77 10.75
E2E Caprese 48.14 13.70 63.37 11.05
CATS 34.42 10.56 61.53 10.11
Layer-wise Caprese 46.32 13.90 61.92 10.82
E2E Caprese 46.17 14.16 63.92 11.03
Gemma 2 9B Instruct 78.17 27.64 63.78 9.91
GRIFFIN 59.21 25.22 63.82 10.14
Layer-wise Caprese 76.72 25.84 64.20 9.82
E2E Caprese 76.65 25.30 64.42 9.92
CATS 76.50 27.32 64.37 9.78
Layer-wise Caprese 77.18 28.16 64.52 10.46
E2E Caprese 77.18 28.00 64.87 10.02
Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). $F$ sparsity is set to 50%, and $r=256$.

### v3

# Scalable LLM Reasoning Acceleration with Low-rank Distillation

@@ Abstract
Due to long generations, large language model (LLM) math reasoning demands significant computational resources and time. While many existing efficient inference methods have been developed with excellent performance preservation on language tasks, they often severely degrade math performance. In this paper, we propose Caprese, a resource-efficient distillation method to recover lost capabilities from deploying efficient inference methods, focused primarily in feedforward blocks. With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the reasoning capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs. Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>16% time-to-next-token reduction) while encouraging response brevity (up to 8.5% fewer tokens).
@@ 1 Introduction
With the increasing capabilities of large language models (LLMs) [52], outputs are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving. These tasks tend to demand long generation which drives up latency, making efficiency a dire issue. Fortunately, many sparsity-based efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization. However, for reasoning tasks, many of these algorithms begin to break down, decimating performance, despite their robustness in language settings. Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and reasoning capabilities.
@@
One of the main differences between reasoning and many language tasks is the generation length. Reasoning typically involves long generations from chain-of-thoughts (CoTs) [53], which often far surpass the length of the input query. However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time. For instance, deploying CATS [31], a sparse thresholding method, on Gemma 2 2B [51] has little impact on language generation performance, but knocks GSM8K [8] accuracy from 51.02% to 34.42% (Table 4). Similarly, for thinking models: applying GRIFFIN [16], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [23] drops MATH-500 [34] accuracy from 79.40% to 42.00% (Table 1). Moreover, since generation is much more computationally demanding than prefill, there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling. Thus, this poses two key inference challenges with math reasoning. First, even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [60]. Second, the already computationally expensive LLM autoregressive decoding process is accentuated with generating CoTs. This problem is further exaggerated as CoTs scale in length in thinking models [23] and as the number of repeated generations scale to elicit higher quality answers [5, 54, 48]. An ideal method should be efficient, performance preserving, and easy to integrate.
@@
Figure 1: A full FF block maximizes accuracy without any benefit to efficiency. Compressed $F$ algorithms can be very efficient by using subsets of the FF block but harm math performance. Our method, Caprese, uses a compressed FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for performative inference in language and math settings while being efficient. Layers are drawn as trapezoids to highlight the expansion of the intermediate feature size compared to the input size.
@@
Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible. (We choose to focus on FF blocks since they contribute around 2/3 of an LLM’s parameters and about 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods. Using an oracle top-$k$ filter on FF nonlinearity output magnitudes, we observe huge reductions in error with a low-rank approximation to the FF output residuals (Figure 2). Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively tiny. This observation motivates us to estimate the residual from FF compression methods with low-rank layers.
@@
We introduce Caprese (CAPability REcovery with Scalable Efficiency) to learn FF residuals from a compressed FF algorithm using small low-rank linear layers for improved performance (Figure 1). While Caprese works for any task, we focus on reasoning (e.g., math) in this paper, as current efficient compression methods can obliterate an LLM’s reasoning capabilities. Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome the aforementioned challenges of reasoning and makes significant progress towards an ideal efficient method:
@@
    Performance Enhancement: Demonstrated across a variety of instruct and thinking models, Caprese recovers much if not all of the reasoning performance lost from deploying a compressed FF algorithm without harming language tasks. Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling.
@@
    Low Budget: The distillation process of Caprese is cheap since we are able to see significant gains in performance by training on only 20K synthetic math samples. Moreover, with a low-rank layer size of only 256, this equates to adding roughly $0.8\%$ additional parameters into Llama 3.1 8B and Gemma 2 9B, dwarfed by savings in active parameters ($\sim$2B cut for both models with Caprese (CATS)).
@@
Section 2 covers related works and describes GRIFFIN and CATS, two efficient sparse FF compression algorithms that serve as baselines. Section 3 details Caprese’s architecture and distillation procedure. Then, we showcase the strong performance of Caprese in Section 4: scaling generation length with thinking LLMs (Section 4.1), scaling generation outputs (Section 4.2), and instruct LLM inference (Section 4.3). Finally, we quantify efficiency improvements in Section 4.4.
@@ 2 Background
Here, we provide an overview of the efficient LLM inference literature that Caprese builds upon with in-depth descriptions of GRIFFIN and CATS which we use as baselines.
@@ 2.1 Related Works
To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [22, 12, 33, 15, 36, 46]. Pruning sets a parameter subset to zero to enforce static sparsity [30, 50, 21, 37]. Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [28, 59, 9, 36]. Similar to MoEs, another line of work aims at exploiting contextual sparsity to dynamically select FF neurons for each input without training. GRIFFIN [16] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation. CATS [31] uses hard thresholding to skip computation in parts of the FF block with a custom kernel. Though not constrained to FF blocks, quantization is another popular method for compression [13, 11, 17, 35].
@@
Several works aim to improve reasoning efficiency by directly reducing the number of generated tokens [45, 41, 3, 4, 55, 42]. Since we aim to reduce the per-token latency, this line of research to encourage brevity is orthogonal to ours and could be used alongside our method.
@@
To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [29, 26]. More recently, there is an increasing effort towards test-time scaling to boost performance on difficult tasks like math. Two ways of test-time scaling that have seen great success. First, for a single prompt, multiple responses can be sampled from the LLM [5, 48, 54, 38, 49, 14]. This increases the probability that a desired response lies in the pool of responses. While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one. Second, CoTs can be lengthened for each response [23, 39, 56]. By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
2.2 Training-free & Adaptively Sparse FF Compression Methods
@@
where $\bm{W}_{g},\bm{W}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$, $\bm{W}_{2}\in\mathbb{R}^{D_{\text{FF}}\times D}$, $\sigma$ is a nonlinear function, $\odot$ is an element-wise multiplication operator, and $D_{\text{FF}}$ is the FF intermediate feature size. Typically, $D_{\text{FF}}\gg D$. Although recent LLMs use this architecture, for LLMs without gated functions (e.g., OPT [58]), $\bm{X}\bm{W}_{1}$ can be removed. Bias terms are omitted for brevity.
@@
GRIFFIN [16] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns. GRIFFIN calculates $\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$. The result $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ gives a metric to perform top-$k$ selection across corresponding columns of $\bm{W}_{g}$ and $\bm{W}_{1}$, and rows of $\bm{W}_{2}$ to produce $\widehat{\bm{W}}_{g},\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{k\times D}$. Then, for the generation phase, the following FF block is used for input $\bm{x}\in\mathbb{R}^{D}$:
@@
CATS [31] uses hard thresholding to skip computation in part of the FF block. Letting $T_{\tau}$ be the hard thresholding function with threshold $\tau$, CATS computes
@@ 3.1 Layer-wise Distillation
Inference efficiency algorithms often introduce feature approximation errors in favor of faster generation, which we mitigate with distillation. Let the efficient and approximate FF block be $\widehat{\text{FF}}(\bm{x})=\widehat{\text{FF}}_{2}(\widehat{\text{FF}}_{1}(\bm{x}))$. In our design of Caprese, we do not constrain $\widehat{\text{FF}}$ to be any specific method or architecture. For instance, $\widehat{\text{FF}}$ could be an FF block with GRIFFIN [16], CATS [31], or quantized weights. Then, the error is
@@
for input set $\mathcal{X}$, $\bm{L}\in\mathbb{R}^{D\times r}$, and $\bm{R}\in\mathbb{R}^{r\times D}$ where $r\ll D_{\text{FF}}$. The optimal solution can be computed analytically since this is a reduced rank regression problem, but the size of $|\mathcal{X}|$ and $D$ may make it prohibitively expensive. Therefore, we opt to learn $\bm{L}$ and $\bm{R}$ independently for every FF block (i.e., previous FF blocks are assumed to be from the original model), allowing for parallel layer-wise training. This takes inspiration from LESS [18] which uses layer-wise training on attention residuals for key-value cache compression. Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen. We also distill end-to-end (E2E) to further improve the performance.
@@ 3.3 Parallel Inference Computation
The computation of $\bm{x}\bm{L}\bm{R}$ can be done in parallel with the original FF operations. In fact, $\bm{L}$ and $\bm{R}$ can be concatenated with the up and down projection matrices, respectively.
In other words, the Caprese FF block is $\widehat{\text{FF}}^{+}(\bm{x})=\widehat{\text{FF}}^{+}_{2}(\widehat{\text{FF}}^{+}_{1}(\bm{x}))$, such that
@@ 3.4 Training Details
We set the inner dimension of our low-rank layer to $r=256$ (to see the effect of different $r$, see Appendix A). In comparison to the large inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B). We use a 20K subset of a common synthetic math training set for training. For a fair comparison, the same subset is used for both layer-wise and E2E distillation for every model. Each training sample is prepended with a CoT instruction: “Please reason step by step.” At test time, the actual instructions may be vastly different. Layer-wise and E2E training consists of 20 epochs and 3 epochs, respectively. Training and inference are done in BF16. Hyperparameters are listed in Appendix G.
@@ 3.5 Comparison with LoRA
Our method derives inspiration from and shares a slight connection with low-rank adaptation (LoRA), a widely used method for efficient fine tuning with the addition of low-rank parameters [27], but remains distinct since they target different inefficiencies. To illustrate, recall a sparse algorithm constructs $\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$ by selecting columns from $\bm{W}_{1}$. Alternatively, construct $\widetilde{\bm{W}}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$ by setting unwanted columns to zero, so nonzero columns in $\widetilde{\bm{W}}_{1}$ match columns in $\widehat{\bm{W}}_{1}$ and vice versa. Doing the same to columns of $\bm{W}_{g}$ and rows of $\bm{W}_{2}$, LoRA learns $\bm{A}_{1},\bm{A}_{g},\bm{B}_{2}^{\top}\in\mathbb{R}^{D\times r}$ and $\bm{B}_{1},\bm{B}_{g},\bm{A}_{2}^{\top}\in\mathbb{R}^{r\times D_{\text{FF}}}$ to form a new FF block with LoRA weights:
@@
Table 1: Thinking models’ 0-shot accuracies on reasoning tasks. Sparsity is set at 50%, and $r=256$. AIME 2024, AMC 2023, and BRUMO 2025 columns also include the sample standard deviation across 4 runs. Further improvements with reselection are shown in Table 3.
Model MATH-500 AIME 2024 AMC 2023 BRUMO 2025 GPQA
DS-Qwen 1.5B 79.40 30.00 $\pm$ 2.72 60.83 $\pm$ 7.47 25.83 $\pm$ 3.19 18.69
GRIFFIN 42.00 2.50 $\pm$ 1.67 16.88 $\pm$ 1.25 1.67 $\pm$ 1.92 11.62
Layer-wise Caprese 47.20 2.50 $\pm$ 1.67 28.13 $\pm$ 2.39 0.00 $\pm$ 0.00 13.13
E2E Caprese 60.40 6.67 $\pm$ 2.72 34.38 $\pm$ 6.57 4.17 $\pm$ 1.67 16.67
CATS 72.00 11.67 $\pm$ 6.94 37.50 $\pm$ 3.54 10.83 $\pm$ 4.19 11.62
Layer-wise Caprese 73.80 15.83 $\pm$ 3.19 48.13 $\pm$ 3.75 13.33 $\pm$ 2.72 16.67
E2E Caprese 74.80 20.83 $\pm$ 1.67 47.50 $\pm$ 3.54 20.00 $\pm$ 2.72 22.22
DS-Qwen 7B 90.20 47.50 $\pm$ 7.87 75.00 $\pm$ 4.08 45.83 $\pm$ 4.19 38.38
GRIFFIN 80.60 21.67 $\pm$ 4.30 62.50 $\pm$ 8.90 25.00 $\pm$ 8.39 21.72
Layer-wise Caprese 84.80 29.17 $\pm$ 3.19 64.38 $\pm$ 7.18 27.50 $\pm$ 4.19 27.78
E2E Caprese 85.40 30.00 $\pm$ 6.09 71.88 $\pm$ 5.54 29.17 $\pm$ 5.69 23.74
CATS 89.20 34.17 $\pm$ 7.88 62.50 $\pm$ 2.04 37.50 $\pm$ 7.39 32.83
Layer-wise Caprese 87.00 35.00 $\pm$ 4.30 70.00 $\pm$ 10.21 39.17 $\pm$ 3.19 38.38
E2E Caprese 90.00 33.33 $\pm$ 4.71 78.13 $\pm$ 5.15 47.50 $\pm$ 5.00 35.86
DS-Qwen 14B 92.80 63.33 $\pm$ 2.72 90.63 $\pm$ 2.39 60.83 $\pm$ 4.19 52.53
GRIFFIN 89.80 32.50 $\pm$ 7.39 80.63 $\pm$ 3.75 33.33 $\pm$ 3.85 41.92
Layer-wise Caprese 90.80 41.67 $\pm$ 5.77 86.88 $\pm$ 4.27 43.33 $\pm$ 4.71 55.05
E2E Caprese 89.20 39.17 $\pm$ 9.18 84.38 $\pm$ 6.57 40.00 $\pm$ 2.72 43.43
CATS 92.80 58.34 $\pm$ 1.92 88.75 $\pm$ 1.44 55.00 $\pm$ 3.33 44.95
Layer-wise Caprese 91.00 61.67 $\pm$ 5.77 88.75 $\pm$ 1.44 53.00 $\pm$ 7.20 50.51
E2E Caprese 92.00 58.34 $\pm$ 9.62 87.50 $\pm$ 2.04 49.17 $\pm$ 4.19 51.01